OpenStudy (anonymous):

Polynomials & Factoring Help!!!!!!!!! 1. What is the degree of the monomial? 3x^10 A) 10 B) 3 C) 13 D) 30 2. What is the degree of the monomial? -2 A) -2 B) -3 C) 0 D) -1 3. What is the sum or difference? 4x^10 - 9x^10 A) -5x^10 B) -5x^20 C) -36x^10 D) -36x^20 4. What is the sum or difference? 7t^5 - 10t^5 A) -17t^5 B) -70t^5 C) -3t^5 D) -3t^10

OpenStudy (michele_laino):

exercise #1 here the degree of the polynomial is given by the exponent of the monomial \(x^{10}\)

OpenStudy (michele_laino):

exerxise #2 here we have no letteral part, and that means all exponents are equal to zero, so what can you conclude?

OpenStudy (michele_laino):

exercise #3 for sum you have to compute \(4-9=...?\) as coefficient of \(x^{10}\) for difference you have to compute \(4-(-9)=...?\)as coefficient of \(x^{10}\)

OpenStudy (michele_laino):

exercise #4 for sum you have to compute \(7-10=...?\) as coefficient of \(t^5\) for difference you have to compute \(7-(-10)=...?\)as coefficient of \(t^5\)